OpenStudy (anonymous):

solve and check solution 1/2+3/t=5/8, i have a few others like this i just don't understand the process

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@Hero @Michele_Laino

OpenStudy (anonymous):

first subtract 1/2 from 5/8

OpenStudy (anonymous):

alright, one moment

OpenStudy (dan815):

okay so first make it common fraction

OpenStudy (anonymous):

wait so i don't do what @10115658 said?

OpenStudy (dan815):

oh lol ya that works too

OpenStudy (dan815):

there are a lot of ways to solve it

OpenStudy (anonymous):

5/8-1/2=1/8

OpenStudy (anonymous):

3/t = 1/8

OpenStudy (anonymous):

Now take the reciprocal of both sides to get T in the numerator

OpenStudy (anonymous):

oh ok ha yea so we're going @10115658's route

OpenStudy (anonymous):

t/3 = 8

OpenStudy (anonymous):

ok gotcha, i understand, is that the final answer ? or is there more?

OpenStudy (anonymous):

how do we get rid of the 1/3?

OpenStudy (anonymous):

multiply both sides by 3

OpenStudy (dan815):

1/3rd of t = 8 so what does that mean t is?

OpenStudy (anonymous):

that will give you the answer

OpenStudy (anonymous):

wait, so responding to what @10115658 said, i multiplied the sides by three and id end up with 3/9 correct? and dan would t be 3.5 ? ha I'm not sure

OpenStudy (anonymous):

>.<

OpenStudy (anonymous):

i hate fractions

OpenStudy (anonymous):